OpenStudy (shamil98):

Dem definite integrals @Luigi0210

OpenStudy (shamil98):

\[\huge \int\limits_{0}^{4} x^2 + 2 dx\]

OpenStudy (shamil98):

first you integrate it..

OpenStudy (shamil98):

\[\huge = \frac{ x^3 }{ 3 } + 2x\]

OpenStudy (shamil98):

then do dis rite? \[\huge \frac{(4)^3 }{ 3 } + 2(4) - ( \frac{ (0)^3 }{ 3 } + 2(0) )\]

OpenStudy (luigi0210):

Correct my good sir

OpenStudy (shamil98):

thank you sir. \[\huge (\frac{ 64 }{ 3 } + 8) - 2 = 88/3 - 2\] \[\huge \frac{ 82 }{ 3 }\]

OpenStudy (luigi0210):

Good job my good sir, you have learned well.

OpenStudy (shamil98):

yay

OpenStudy (luigi0210):

And btw, if you have a calculator, you can do it on there too :D

OpenStudy (shamil98):

yeah, on a graphing calc, i have one but have no idea how to use it lol

OpenStudy (luigi0210):

I can teach you, but what kind of graphing calc?

OpenStudy (shamil98):

ti-84 plus

OpenStudy (luigi0210):

Good, we got dis then ლ(́◉◞౪◟◉‵ლ)

OpenStudy (luigi0210):

Okay, so let's try it mathematically first: \[\LARGE \int\limits_{0}^{10}~2x~dx\] Solve this simple one

OpenStudy (shamil98):

\[\huge \frac{ 2x^2 }{ 2 }\] \[\huge x^2\] do i have to write the constant C in this though?

OpenStudy (shamil98):

\[\huge (10)^2 - (0)^2 = 100\]

OpenStudy (luigi0210):

Nope, +/-C is only in indefinite integrals. And correct, now get your calculator

OpenStudy (shamil98):

Oh, i see. got it with me.

OpenStudy (luigi0210):

So start off by graphing 2x

OpenStudy (shamil98):

done

OpenStudy (luigi0210):

Now that you're on the graphing screen, press "2nd" and then to the top press the button that says "Trace"

OpenStudy (shamil98):

and option 7: it has the integral sign and f(x) dx that one?

OpenStudy (luigi0210):

Correct, so press that one

OpenStudy (luigi0210):

And it will give you a "Lower Limit?" and "Upper Limit" which you can just type in with the keypad thing.

OpenStudy (luigi0210):

Those limits are the ones on the integral

OpenStudy (luigi0210):

So it should be Lower Limit 0, enter, and Upper Limit 10, enter, and you should get 100

OpenStudy (shamil98):

yeah i got 100

OpenStudy (shamil98):

sweet, thanks :D

OpenStudy (luigi0210):

Anytime brah